# Effective (Floquet) Lindblad generators from spectral unwinding

### v1

@@ Abstract
where $\Delta_{\ell}$ denotes the level splitting for the $\ell$-th spin and $w$ is the coupling strength between neighbouring spins. The driving is parameterized by the driving strength $E$ and -frequency $\omega$. As we show in Fig. 1(a) and (c), even if we calculate the dynamical map $\mathcal{V}(T)$ for an evolution with time-dependant von-Neumann generator $\mathcal{G}(t)=-i[H(t),\cdot]$, the generator $\mathcal{S}_{0}$ of the principle branch is not necessarily of von-Neumann form $\mathcal{N}$ (and may even be not of Lindblad form, cf. blue areas). This is in stark contrast to our knowledge of the existence of the Floquet Hamiltonian $H_{\mathrm{F}}$, i.e. $\mathcal{S}=-i[H_{\mathrm{F}},\cdot]$ is a valid effective Lindblad generator. Only after also searching around the first two neighbouring branches, i.e. $x_{c}\in\{-1,0,1\}$, one finds again a valid von-Neumann generator, as we show in Fig. 1(b). However, now, for every point in the phase diagram, we have a maximal number of $3^{2^{2L-1}-1}$ branches for which condition (ii) has to be checked. In the case of a spin chain of length $L=2$ this can still be computed in a reasonable time. However, due to the super exponential scaling, for $L=3$ ($L=4$) we would already have a maximal number of $\approx 6\cdot 10^{14}$ ($\approx 3\cdot 10^{60}$) branches to consider, simply to avoid the winding problem (cf. yellow line in Fig. 2(a)). Thus, in order to avoid this problem, we will now propose a different approach, which is based on a strategy for unwinding the quasienergy spectrum.
@@
Spectral unwinding.—For undriven systems it is straight forward to define an “unwinding” procedure that undoes this winding into the first Floquet Brioullin zone: Since the Floquet modes ${|{\Phi_{c}(t)}\rangle\!\rangle}$ with complex quasienergy differences $\Omega_{c}$ are time-periodic, we can expand them in a Fourier series ${|{\Phi_{c}(t)}\rangle\!\rangle}=\sum_{n\in\mathbb{Z}}e^{i\omega nt}{|{\Phi_{c}^{(n)}}\rangle\!\rangle},$ and compute the trace norm for all of their Fourier components $||{|{\Phi_{c}^{(n)}}\rangle\!\rangle}||=\operatorname{Tr}\sqrt{\Phi_{c}^{\dagger(n)}\Phi_{c}^{(n)}}$. For every mode $c$ we determine $x_{c}^{(\text{max})}=\underset{x\in\mathbb{Z}}{\arg\max}\{||{|{\Phi_{c}^{(x)}}\rangle\!\rangle}||\}.$ Then, we shift the quasienergies $\Omega_{c}$ by this number of quanta $x_{c}^{\text{(max)}}$ which gives the ‘unwound’ Floquet generator $\hat{\mathcal{S}}_{\text{unf}}=\hat{\mathcal{S}}_{0}+i\frac{2\pi}{T}\sum_{c=1}^{N_{c}}x_{c}^{(\text{max})}\bigl{(}{|{\Phi_{c}}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{c}}|}-\mathbb{F}{|{\Phi_{c}}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{c}}|}\mathbb{F}\bigr{)}.$ In the undriven system, this gauge transformation removes the micromotion and transforms the Floquet modes into the static eigenmodes of the time-independent generator $\mathcal{G}$. In the following, we show that this procedure also significantly improves the search for an effective generator in the case of a time-dependent generator $\mathcal{G}(t)$. In this regard, we go back to our spin chain model. In Fig. 1(d) and (f) we show the phase diagrams we obtain for the generator of the principle branch $\mathcal{S}_{0}$ and the unwound generator $\mathcal{S}_{\text{unf}}$ for chain lengths $L=2$ and $L=3$, respectively. Here, between $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$, we always choose the generator with the smallest distance from Markovianity and plot its distance $\mu$. In the Supplemental Material (SM) [59], we plot the phase diagrams of $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$ seperately to show how they complement each other. By using the procedure above, in comparison to Fig. 1(a) and (c), where we have only considered $\mathcal{S}_{0}$, we find that specifically in the high-frequency regime we obtain a lot more points $(E,\omega)$ give generators of von-Neumann form. (In the SM [59], we introduce a measure to determine the distance from the von-Neumann form and show that both measures almost coincide for our problem.)
@@
However, as we observe in Fig. 1(d),(f) for low frequencies, this method of unwinding fails. This is expected, since in this regime the peaks in the Fourier spectra can become more broadly distributed over several Fourier modes and indistinctly peaked. To overcome this issue, one has to include additional frequencies for the search in this regime. In the following, we outline how to find well-chosen candidates for effective generators to test for Markovianity, other than $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$. To this end, we set $x_{c}^{(\text{max,0})}=x_{c}^{(\text{max})}$ and chose a number $\eta\in[0,1]$, which determines at which ratio between the amplitudes of the neighbouring peaks at $x_{c}^{(\mathrm{max},i>0)}$ and the main peak at $x_{c}^{(\text{max,0})}$ we want to include them into the generation of branch combinations. The family of generators we consider therefore is $\hat{\mathcal{S}}_{\{\vec{\xi}\}}$ with the branches according to Eq. (3) but here, the integers $\xi_{c}$ are not any arbitrary integers, but the frequencies of the main and the included neighbouring peaks in the Fourier spectra of every Floquet mode
@@
where for every pair of complex modes $c$, we determine the number $z_{c}\in\mathbb{N}$ which counts how many neighbouring peaks have amplitudes that are on the order of the amplitude of the main peak according to the criterion ${A^{(\text{max,i})}_{c}}/{A^{(\text{max,0})}_{c}}\geq\eta,$ with $A^{(\text{max},i)}_{c}=\operatorname{Tr}\sqrt{\Phi_{c}^{\dagger(x^{(\text{max,i})})}\Phi_{c}^{(x^{(\text{max,i})})}}$. We additionally introduce a cutoff $N_{b}$ for the maximal number of branches considered and only consider the peaks $i=1,...,N_{b}$. This gives us a way to systematically search around the unwound generator $\mathcal{S}_{\text{unf}}$, instead of searching around the generator of the principle branch $\mathcal{S}_{0}$, which can drastically reduce the required numerical effort. In Fig. 1(e) we apply this modified Markovianity test for a spin chain of length $L=2$ and the parameters $N_{b}=2$ and $\eta=0.7$. We observe that we almost fully recover valid von-Neumann generators. Note that for $L\geq 3$ this choice of $N_{b}$ and $\eta$ will not suffice to recover valid von-Neumann generators in every point of the phase diagram. Here, to minimise the complexity, for a given point one can search for the smallest $N_{b}$ and largest $\eta$ which yield a generator of von-Neumann form. In the SM [59], for $L=3$ we have set $N_{b}=1$ and varied $\eta$ to find an optimum along the vertical red line at $E=2.5$ in Fig. 1(e).
@@
To compare the reduction in complexity of our method to the naive procedure of searching around the principle branch, in Fig. 2(a) we plot the cardinality $|M|$ of the set $M$ of all possible branch combinations for which we test for Markovianity. For the standard Markovianity test the size of $M$ is dependent on the included neighboring branches $x_{c}$. For the modified Markovianity test the size of $M$ depends on our choice of $N_{b}$ and $\eta$, which we choose such that we minimise the included branches but still get a proper von-Neumann generator. For the principle branch $\mathcal{S}_{0}$ we always have $|M|=1$ (blue line). The generators of the principle branch yields poor results due to the winding of quasienergy differences. Including the first two neighbouring, i.e. $x_{c}\in\{-1,0,1\}$, fixes this problem, but greatly increases the complexity as we see for the orange line. On the other hand, trying to revert the winding by only including $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unf}}$, we have $|M|=2$ (green line). Using the modified Markovianity test of Eq. (8), the minimal cardinality needed to obtain a fully von-Neumann phase diagram for $L=2$ and to make the points on the red line in Fig. 1(c) fully von-Neumann for $L=3$, is plotted by the red line in Fig. 2(a). For a spin chain of length $L=3$ we have $|M|\approx 10^{5}$ on average (see SM [59] for details) to make the red line fully von-Neumann. Searching around the principle branch requires $|M|\approx 10^{13}$. This shows that searching around the unwound generator dramatically decreases the complexity (as long as we are not at low frequencies). Nevertheless, even for the modified method, the complexity still grows fast with the dimensionality of the problem, which is why for bigger systems one might only be able to include the generator of the principle branch $\mathcal{S}_{0}$ and the unwound generator $\mathcal{S}_{\text{unf}}$ (or alternatively, combine our method with the existing machine learning method of Ref. [57]).
@@
Dissipative case.—In the case of a coherent time evolution we can always compute a Floquet Hamiltonian $H_{\mathrm{F}}$ and then obtain the effective time-independent Floquet generator $\mathcal{S}=-i[H_{\mathrm{F}},\cdot]$, which always has von-Neumann form. Such a bypass, however, is not possible in the case of dissipative systems where we search for effective time-independent generators of Lindblad form, i.e. Floquet Lindbladians [41, 42], since here we have to compute $\log(\mathcal{V}(T))/T$ and test these generators for Markovianity [41]. If we do not change the Hamiltonian but add very weak dissipation, we can see exactly the same non-Markovian areas arising in the principle branch (due to the winding of the eigenvalues of $\log[\mathcal{V}(T)]/T$), cf. Fig. 2(b). Here, we add dissipation described by Lindblad jump operators for the quantum-optical master equation $L^{\ell}_{kq}=\sqrt{R^{\ell}_{kq}}{|{\psi_{k}}\rangle}{\langle{\psi_{q}}|},$ where, for simplicity, we neglect the impact of the driving term on the dissipator. Hence, ${|{\psi_{k}}\rangle},E_{k}$ are the many-body eigenstates and -energies of the undriven Hamiltonian for $E=0$. The corresponding jump rate for contact with a reservoir at temperature $\mathcal{T}$ at site $l$ reads $R^{\ell}_{kq}=2\pi\gamma^{2}|{\langle{\psi_{k}}|}\sigma_{x}^{\ell}{|{\psi_{q}}\rangle}|^{2}g(E_{k}-E_{q})$ with $g(E)=E/(e^{E/\mathcal{T}}-1)$, assuming an ohmic bath [60]. By summing over all jump operators we obtain the total time-periodic Lindbladian $\mathcal{L}(t)[\cdot]=-{i}\bigl{[}H(t),\cdot\bigr{]}+\sum_{\ell=1}^{L}\sum_{k,q}\bigl{(}L^{\ell}_{kq}\cdot L_{kq}^{\ell\dagger}-\frac{1}{2}\bigl{\{}L_{kq}^{\ell\dagger}L^{\ell}_{kq},\cdot\bigr{\}}\bigr{)}.$ Only considering the generator of the principle branch $\mathcal{S}_{0}$, for $L=3$ and $L=4$, we obtain the map shown in Fig. 2(b) and (c), respectively. Applying the Markovianity test to the unwound Floquet generator $\mathcal{S}_{\text{unf}}$ and $\mathcal{S}_{0}$ yields the phase diagrams in Fig. 2(d) and (e), where we observe that the phase diagrams contain more valid Floquet Lindbladians $\mathcal{L}_{\mathrm{F}}$ (white region) [41, 42] or have significantly lower distances from Markovianity $\mu$ in the areas where no $\mathcal{L}_{\mathrm{F}}$ exists. Note again that scanning through all the possible logarithm branches would mean to scan through possibly $\approx 10^{14}$ candidates for $L=3$ and $\approx 10^{60}$ candidates for $L=4$, if we merely wanted to include the first two neighbouring branches $x_{c}\in\{-1,0,1\}$. The unwound generator $\mathcal{S}_{\text{unf}}$ thus gives rise to a fast way of determining a particularly good candidate generator which we can test for Markovianity, when dealing with high-dimensional open quantum many-body systems. In the Supplemental Material we give an example to show that the unwound generator is a good candidate not only for Floquet systems but also in the general case of a time-dependent non-periodic map $\mathcal{V}(t)$. We show that for a convex combination $\mathcal{V}(t)=\lambda\mathcal{V}_{1}(t)+(1-\lambda)\mathcal{V}_{2}(t)$ [61], where $\mathcal{V}_{j}(t)=e^{t\mathcal{L}_{j}}$ and $\lambda\in[0,1]$, the unwound generator $\mathcal{S}_{\text{unf}}$ produces valid effective Lindbladians in areas where the generator of principle branch $\mathcal{S}_{0}$ does not.

### v3

@@ III Systems with access to full time-dependent map
with level splitting $\Delta_{\ell}$, coupling strength $w$ between neighbouring spins, driving strength $E$ and -frequency $\omega$. As we show in Fig. 1(a) and (c), even if we calculate the dynamical map $\mathcal{V}(T)$ for an evolution with time-dependant von-Neumann generator $\mathcal{G}(t)=-i[H(t),\cdot]$, the generator $\mathcal{S}_{0}$ of the principle branch is not necessarily of von-Neumann form $\mathcal{N}$ (and neither of Lindblad form), see blue areas. This is in stark contrast to our knowledge of the existence of the Floquet Hamiltonian $H_{\mathrm{F}}$, i.e. $\mathcal{S}=-i[H_{\mathrm{F}},\cdot]$ is a valid effective Lindblad generator. Only after also searching around the first two neighbouring branches, i.e. $x_{c}\in\{-1,0,1\}$, one finds again a valid von Neumann generator, as we show in Fig. 1(b). However, for every point in the phase diagram, a maximal number of $3^{2^{2L-1}-1}$ branches has to be checked for condition (ii). Due to the super exponential scaling, for $L=3$ ($L=4$) we would already have a maximal number of $\approx 6\cdot 10^{14}$ ($\approx 3\cdot 10^{60}$) branches to consider, simply to avoid the winding problem (cf. yellow line in Fig. 2(a)). In order to avoid this problem, we propose a different approach which is based on a strategy for unwinding the quasienergy spectrum.
IV Spectral unwinding
For undriven systems it is straightforward to define an “unwinding” procedure that undoes this winding into the first Floquet Brioullin zone: Since the Floquet modes ${|{\Phi_{c}(t)})}$ with corresponding complex quasienergy differences $\Omega_{c}$ are time-periodic, we can expand them in a Fourier series
${|{\Phi_{c}(t)})}=\sum_{n\in\mathbb{Z}}e^{i\omega nt}{|{\Phi_{c}^{(n)}})},$ (14)
and compute the trace norm for all of their Fourier components $||{|{\Phi_{c}^{(n)}})}||=\operatorname{Tr}\sqrt{\Phi_{c}^{\dagger(n)}\Phi_{c}^{(n)}}$. For every mode $c$ we determine
$x_{c}^{(\text{max})}={\arg\max}_{x\in\mathbb{Z}}\{||{|{\Phi_{c}^{(x)}})}||\}.$ (15)
Then, we shift the $\Omega_{c}$ by this number of quanta $x_{c}^{\text{(max)}}$ giving the ‘unwound’ Floquet generator
$\begin{split}\hat{\mathcal{S}}_{\text{unw}}=\hat{\mathcal{S}}_{0}+i\omega\sum_{c=1}^{N_{c}}x_{c}^{(\text{max})}\Bigl{[}{|{\Phi_{c}(0)})}{({\tilde{\Phi}_{c}(0)}|}\\ -\mathbb{F}{|{\Phi_{c}(0)})}{({\tilde{\Phi}_{c}(0)}|}\mathbb{F}\Bigr{]}\end{split}$ (16)
@@ IV Spectral unwinding
In the undriven system, this gauge transformation removes the micromotion and transforms the Floquet modes into the static eigenmodes of the time-independent generator $\mathcal{G}$ [20, 61]. We now show that this procedure also significantly improves the search for an effective generator in the case of a time-dependent generator $\mathcal{G}(t)$. In Fig. 1(d) and (f) we show the phase diagrams we obtain from checking the principle branch $\mathcal{S}_{0}$ and the unwound generator $\mathcal{S}_{\text{unw}}$ for chain lengths $L=2$ and $L=3$. Here, between $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unw}}$, we choose the generator with shortest distance from Markovianity $\mu$. In Appendix A, we plot the phase diagrams of $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unw}}$ separately. By using the procedure above, in comparison to Fig. 1(a) and (c), we find that specifically in the high-frequency regime we obtain a lot more points $(E,\omega)$ yielding generators of von Neumann form. (In App. B, we introduce a measure to determine the distance from the von-Neumann form and show that both measures almost coincide for our problem.)
@@
However, as we observe in Fig. 1(d),(f) for low frequencies, this method of unwinding fails. This is expected, since in this regime the peaks in the Fourier spectra can become more broadly distributed over several Fourier modes and indistinctly peaked, which we illustrate in App. C, by specifically analyzing the Fourier spectra of $||{|{\Phi_{c}^{(n)}})}||$ at points, where $\mathcal{S}_{\text{unw}}$ fails and succeeds in giving a valid generator of von Neumann form. To overcome this issue, one has to include additional frequency peaks for the search in this regime. In the following, we outline how to find well-chosen candidates for effective generators to test for Markovianity, other than $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unw}}$. To this end, we set $x_{c}^{(\text{max,0})}=x_{c}^{(\text{max})}$ and choose a number $\eta\in[0,1]$, which determines at which amplitude ratio we want to include the neighbouring peaks $x_{c}^{(\mathrm{max},i>0)}$ into the possible branch combinations. The resulting family of generators $\hat{\mathcal{S}}_{\{\vec{\xi}\}}$ we consider is given by
@@
where for every pair of complex modes $c$, we include the $z_{c}$ neighbouring Fourier peaks with amplitudes ${A^{(\text{max,i})}_{c}}/{A^{(\text{max,0})}_{c}}\geq\eta,$ with $A^{(\text{max},i)}_{c}=||{|{\Phi_{c}^{(x^{(\text{max,i})})}})}||$. We additionally introduce a cutoff $N_{b}$ for the maximal number of branches considered, i.e. $z_{c}\leq N_{b}$. This gives rise to a systematic search around the unwound generator $\mathcal{S}_{\text{unw}}$, instead of searching around the generator of the principle branch $\mathcal{S}_{0}$, which can drastically reduce the required numerical effort. In Fig. 1(e) we apply this modified Markovianity test for a spin chain of length $L=2$ and parameters $N_{b}=2,\eta=0.7$. We observe that we almost fully recover valid von Neumann generators. Note that for $L\geq 3$ this choice of $N_{b}$ and $\eta$ will not suffice to recover valid von Neumann generators in every point of the phase diagram. Here, to minimise the numerical effort, for a given point, one can search for the smallest $N_{b}$ and largest $\eta$ which yield a generator of von Neumann form. In App. D, for $L=3$ we have set $N_{b}=1$ and varied $\eta$ to find an optimum along the vertical red line at $E=2.5$ in Fig. 1(e).
@@
To compare the reduction in complexity of our method to the naive procedure of searching around the principle branch, in Fig. 2(a) we plot the cardinality $|M|$ of the set $M$ of all possible branch combinations for which we test for Markovianity. For the principle branch $\mathcal{S}_{0}$ we always have $|M|=1$ (blue line). The generators of the principle branch yield poor results due to the winding of quasienergy differences. Including the first two neighbouring, i.e. $x_{c}\in\{-1,0,1\}$, fixes this problem, but greatly increases the complexity as we see for the orange line. On the other hand, trying to revert the winding by only including $\mathcal{S}_{0}$ and $\mathcal{S}_{\text{unw}}$, we have $|M|=2$ (green line). Using the modified Markovianity test of Eq. (17), the red line in Fig. 2(a) shows the minimal cardinality needed to obtain a fully von Neumann phase diagram for $L=2$, as well as to make the points on the red line in Fig. 1(c) fully von Neumann for $L=3$. For $L=3$ we have $|M|\approx 10^{5}$ on average (see App. D for details). Searching around the principle branch requires $|M|\approx 10^{13}$. This shows that searching around the unwound generator dramatically decreases the complexity (as long as we are not at low frequencies, where the peaks in the Fourier spectra may spread across multiple Fourier modes). For systems with large Hilbert space dimensions, where it is numerically impossible to test all branch combinations when considering e.g. the principal and first two neighboring branches, i.e. $x_{c}\in\{-1,0,1\}$, the conventional method would require to blindly guess a branch combination with no physical or mathematical justification, while our method provides a systematic way of searching for new generators in cases where the principle branch $\mathcal{S}_{0}$ and unwound generator $\mathcal{S}_{\text{unw}}$ yield poor results. Nevertheless, even for the modified method, the complexity still grows fast with the dimensionality of the problem, which is why for bigger systems one might only be able to include the principle branch $\mathcal{S}_{0}$ and the unwound generator $\mathcal{S}_{\text{unw}}$ (or alternatively, combine our method with the existing machine learning method of Ref. [58]).
V Dissipative case
In the case of a coherent time evolution we can always compute a Floquet Hamiltonian $H_{\mathrm{F}}$ and then obtain the effective time-independent Floquet generator $\mathcal{S}=-i[H_{\mathrm{F}},\cdot]$, which always has von Neumann form. Such a bypass, however, is not possible in the case of dissipative systems where we search for effective time-independent generators of Lindblad form, i.e. Floquet Lindbladians [41, 42]. Here we have to compute $\log(\mathcal{V}(T))/T$ and test these generators for Markovianity [41]. If we do not change the Hamiltonian but add very weak dissipation, we can see exactly the same non-Markovian areas arising in the principle branch, cf. Fig. 2(b). Here, we add dissipation described by Lindblad jump operators for the quantum-optical master equation
$L^{\ell}_{kq}=\sqrt{R^{\ell}_{kq}}{|{\psi_{k}}\rangle}{\langle{\psi_{q}}|},$ (18)
where, for simplicity, we neglect the impact of the driving term on the dissipator. Hence, ${|{\psi_{k}}\rangle},E_{k}$ are the many-body eigenstates and -energies of the undriven Hamiltonian for $E=0$, respectively. The corresponding jump rate for contact with a reservoir at temperature $\mathcal{T}$ at site $l$ reads
$R^{\ell}_{kq}=2\pi\gamma^{2}|{\langle{\psi_{k}}|}\sigma_{x}^{\ell}{|{\psi_{q}}\rangle}|^{2}g(E_{k}-E_{q})$ (19)
with $g(E)=E/(e^{E/\mathcal{T}}-1)$, assuming an ohmic bath [62].
By summing over all jump operators we obtain the total time-periodic Lindbladian
$\mathcal{L}(t)[\cdot]=-{i}\bigl{[}H(t),\cdot\bigr{]}+\sum_{\ell=1}^{L}\sum_{k,q}\bigl{(}L^{\ell}_{kq}\cdot L_{kq}^{\ell\dagger}-\frac{1}{2}\bigl{\{}L_{kq}^{\ell\dagger}L^{\ell}_{kq},\cdot\bigr{\}}\bigr{)}.$ (20)
@@ V Dissipative case
Only considering the generator of the principle branch $\mathcal{S}_{0}$, for $L=3$ and $L=4$, we obtain the maps shown in Fig. 2(b) and (c), respectively. Applying the Markovianity test to the unwound Floquet generator $\mathcal{S}_{\text{unw}}$ and $\mathcal{S}_{0}$ yields the phase diagrams in Fig. 2(d) and (e), where we observe that the phase diagrams contain more valid Floquet Lindbladians $\mathcal{L}_{\mathrm{F}}$ (white region) [41, 42] or have significantly lower distances from Markovianity $\mu$ in the areas where no $\mathcal{L}_{\mathrm{F}}$ exists. Note again that scanning through all the possible logarithm branches would mean to scan through possibly $\approx 10^{14}$ candidates for $L=3$ and $\approx 10^{60}$ candidates for $L=4$, if we merely wanted to include the first two neighbouring branches $x_{c}\in\{-1,0,1\}$. The unwound generator $\mathcal{S}_{\text{unw}}$ thus gives rise to a fast way of determining a particularly good candidate generator which can be tested for Markovianity, when dealing with high-dimensional open quantum many-body systems. In App. E we provide further observations for interacting systems in terms of the existence of the Floquet Lindbladian. Based on that, we motivate our specific choice of dissipation in this paper. In App. F we further give an example to show that the unwound generator is a good candidate not only for Floquet systems but also in the general case of a time-dependent non-periodic map $\mathcal{V}(t)$. We show that for a convex combination
$\mathcal{V}(t)=\lambda\mathcal{V}_{1}(t)+(1-\lambda)\mathcal{V}_{2}(t),$ (21)
where $\mathcal{V}_{j}(t)=e^{t\mathcal{L}_{j}}$ with Lindbladians $\mathcal{L}_{j}$ and $\lambda\in[0,1]$ (which previously has been studied in the context of definitions of Markovianity [63]) the unwound generator $\mathcal{S}_{\text{unw}}$ produces valid effective Lindbladians in areas where the generator of principle branch $\mathcal{S}_{0}$ does not.
VI Summary